If g (x)=7x-a where a is a constant and g(10)=3g(20) then g(30)=
-175
175
70
35

Answer by ikleyn(52781)

.
1.  Your first step is to determine the value of "a", the constant term.


    For it, use the condition g(10) = 3*g(20).


    After substituting the values to the formula, it will give you an equation


        7*10 - a = 3*(7*20-a).


    Simplify it and find "a".



2.  Having the value of "a", calculate g(30).


Good luck and happy learning / solving / calculating !